# Long-Lived Electronic Coherences from First Principles

### v1

@@ I Introduction
Electronic coherences can be leveraged to control light-driven dynamics in both naturalDuan et al. (2017); Maiuri et al. (2018) and syntheticMerritt et al. (2021); Lépine et al. (2014) molecules. In an isolated atom, coherences can survive for hundreds of picoseconds or more.Tenwolde et al. (1989) Yet in molecules with more than a few atoms, the loss of electronic coherence can be ultrafast, with the vibrations of the molecule acting as a bath that kills coherence.Schwartz et al. (1996) With an eye towards control, it would be extremely desirable to understand the factors determining the rate of electronic decoherence in excited molecules.
@@
One can clearly see that a description of the vibrational wave packets on multiple electronic states is central to a correct quantitative description of electronic coherence and decoherence. Thus methods based on a fully quantum treatment of nuclear dynamics are a natural fit, providing an accurate description of coherence, albeit at a high computational cost. For example, multi-configuration time-dependent Hartree (MCTDH)Csehi et al. (2020); Dey et al. (2022); Arnold et al. (2017) can provide an extremely accurate description of the dynamics of vibrational wave packets on multiple electronic states, but requires a fitted PES, often in reduced dimensionality to keep computational costs down. On-the-fly ab initio methods based on Gaussian wave packetsVacher et al. (2017); Jenkins et al. (2018); Tran et al. (2020); Scheidegger et al. (2022) provide a lower cost description in full dimensionality. But they come with their own complexities, owing to the need for multiple trajectory basis functions evolving on different PESs to interact with one another. Classical Ehrenfest simulations can provide useful insights into coherences on short timescales,Jenkins et al. (2016); Franco and Brumer (2012) but Ehrenfest does not incorporate decoherence and, therefore, cannot discern between long- and short-lived coherences. This leads to well-known qualitative errors on longer timescales.Tully (1990); Cotton and Miller (2013) Incorporating decoherence effects into Ehrenfest effectively solves the problem for two-state systems,Zhu et al. (2004); Bedard-Hearn et al. (2005) but properly treating long-lived coherences in systems with many electronic states is not straightforward.Esch and Levine (2020a) Thus, an on-the-fly classical trajectory-based approach capable of accurately describing both long-lived coherences and ultrafast decoherence has been elusive.
@@
To fill this gap, we have introduced the Ehrenfest with collapse-to-a-block (TAB) nonadiabatic molecular dynamics approach,Esch and Levine (2020a, 2021) and in this work we apply it to simulate the long-lived coherence observed in excited thiophene. The TAB method has recently been combined with an ab initio treatment of the electronic structure for the first time.Suchan et al. (2024) In the present work, we demonstrate TAB to be a computationally lightweight alternative to fully quantum methods, capable of discerning between long- and short-loved coherences. Ab initio TAB simulates the system on the fly in its full dimensionality and does not require fitted PESs. Used in conjunction with time-dependent configuration interaction (TDCI)Peng et al. (2018); Durden and Levine (2022) calculations of multiphoton excitation by the pump pulse, we are able to explicitly simulate the dynamics observed in the Kaufman experiment. With detailed simulations of coupled electron-nuclear dynamics in hand, we definitively assign the observed coherences to specific states and gain understanding of the physics underlying its long lifetime.
@@ II.1 Nonadiabatic dynamics: TAB method
A detailed description of the TAB algorithm can be found in Refs. 22 and 27. Here, we highlight the key features of the method relevant to the present results. The TAB method is based on the Ehrenfest molecular dynamics method,Miller and McCurdy (1978); Meyera) and Miller (1979); Micha (1983); Isborn et al. (2007) where the nuclei are treated classically, and the force vector is calculated in a mean field fashion:
$\mathbf{F}(t)=-\bra{\psi(t)}\nabla_{\bm{R}}\hat{H}_{\mathrm{e}}(\bm{R}(t))\ket{\psi(t)}_{\bm{r}}\,.$ (5)
The electronic wave function is then evolved by solving the time-dependent electronic Schrödinger equation:
@@
In the Ehrenfest method, there is no consideration of decoherence due to the loss of nuclear overlap (mechanism c above).Kirrander and Vacher (2020) As noted above, this leads to the evolution of classical trajectories along the mean-field PES, which can result in unphysical outcomes.Tully (1990); Cotton and Miller (2013) We remedy these issues by incorporating the collapse-to-a-block (TAB) decoherence correction scheme.Esch and Levine (2020a, 2021, b)
The specific version of the TAB collapse algorithm used here is described in detail ref 27, where this correction is referred to as TAB-w1. The TAB correction assumes Gaussian decay of pairwise electronic coherences with a rate governed by the difference of the gradients of associated PESs, as originally derived by Rossky and Bittner:
@@
where $\eta$ indexes vibrational degrees of freedom, $F_{i,\eta}$ is the corresponding force matrix element on state $i$, and $\alpha_{\eta}$ is a parameter that represents the width of a Gaussian vibrational wave packet. This expression has a very intuitive interpretation; the more different the slopes of the PESs of two states, the faster their vibrational overlap, $\chi^{\prime}_{1}(\bm{R},t)\chi^{\prime*}_{2}(\bm{R},t)$, will decay. This and similar approximations have been used for years to correct the nonadiabatic transition probabilities incorporated into simulations of nonadiabatic molecular dynamics.Zhu et al. (2004); Bedard-Hearn et al. (2005); Ouyang et al. (2015); Min et al. (2015); Subotnik et al. (2016); Wang et al. (2016); Martens (2019); Vindel-Zandbergen et al. (2021) The novel feature of TAB is that each state pair is assigned an independent decoherence time, allowing for an accurate treatment of decoherence in systems with more than two electronic states.Esch and Levine (2020a)
@@
A complete definition of $w_{ij}(\tau_{ij},\dot{\rho}_{ii},\dot{\rho}_{jj})$ may be found in ref 27. The diagonal elements of the target density matrix are not scaled,
@@
Such blocks might be composed of a single state or any combination of available states while retaining the population ratio and the phase angles between all pairs. We use the algorithms described in Ref. 27 to determine the set $\{\bm{\rho}_{a}^{\mathrm{block}}\}$ and the corresponding probabilities, $\{P_{a}^{\mathrm{block}}\}$. The sum of these probabilities must be unity, and a constrained linear least squares solution to Eq. 10 ensures that all probabilities are non-negative.
Finally, we stochastically decide whether the wavefunction will collapse into one of pure states represented by these blocks. A random number between 0 and 1 is selected and compared to the probabilities, $\{P_{a}^{\mathrm{block}}\}$, to determine which state the electronic subsystem collapses into. After collapse, the momentum is scaled to ensure conservation of energy. This block-wise algorithm allows TAB to correctly describe the decay of coherences between all pairs of states—even when many states are present, pairs of states with parallel PESs maintain coherence indefinitely, while very non-parallel pairs lose coherence rapidly.Esch and Levine (2020a)
In this work, we use a fully ab initio method to propagate the electronic wave function: time-dependent configuration interaction.Peng et al. (2018) In TDCI, the wave function is expressed as a linear combination of Slater determinants, $\Phi_{I}$, with time-dependent expansion coefficients $C_{I}(t)$,
@@
$i\dot{\bf{C}}(t)={\bf{H}_{\mathrm{e}}}(t){\bf{C}}(t)$ (12)
and solve the equation numerically using the symplectic split operator integrator.Blanes et al. (2006) In this work, we select a complete active space (TD-CASCI) ansatz. Our GPU-accelerate implementation is in the TeraChem electronic structure software package.Seritan et al. (2021, 2020); Durden and Levine (2022); Fales and Levine (2015); Peng et al. (2018); Hohenstein et al. (2015)
We have previously applied ab initio TAB to predict the ultrafast electronic diffraction spectrum of the photodissociation of cyclobutanoneSuchan et al. (2024) in the context of a community prediction challenge.Miao et al. (2024); Martín Santa Daría et al. (2024) Our prediction, though slightly too fast, is in good agreement with experiment. [^1] Preliminary experimental data communicated by Thomas Wolf. However, that problem involved a small number of electronic states and only short-lived coherences. The present work represents the first application of TAB to describe long-lived electronic coherences in a system with many electronic states, the class of problems for which it is almost uniquely well suited.
@@ II.2 Thiophene’s electronic structure
Despite its apparent simplicity, previous theoretical studies reveal that accurately describing the electronic excited states of thiophene (C4H4S) can be quite challenging. The HOMO ($\pi_{1}$) and HOMO-1 ($\pi_{2}$) orbitals, from which the low-lying excitations originate, have $\pi$ character. The excitations lead not only to antibonding valence orbitals ($\pi^{*}$ and $\sigma^{*}$), but also to a set of low-lying Rydberg orbitals. This puts additional demands on the atomic basis set and wave function expansion. The difference in electronic correlation energy between the valence and Rydberg states is quite large, as demonstrated in many past studies based on configuration interactionPalmer et al. (1999), complete active space self-consistent field (CASSCF)Serrano-Andrés et al. (1993); Heller et al. (2021); Pastore et al. (2007), multireference perturbation theory Kaufman et al. (2022); Parkes and Worth (2024), coupled cluster theory Pastore et al. (2007); Holland et al. (2014), and other methods.Huang et al. (2008); Salzmann et al. (2008); Kleinschmidt et al. (2002); Jones et al. (2019)
For computational feasibility, we must choose a relatively inexpensive electronic structure method for use in conjunction with TAB. In addition, TAB makes use of a time-dependent electronic structure methods (i.e. one that solves the time-dependent electronic Schödinger equation) unlike typical ab initio molecular dynamics methods, which use a standard time-independent approach. To this end, we choose the time-dependent (TD-) floating occupation molecular orbitals (FOMO-) complete active space configuration interaction (CASCI) approach.Granucci et al. (2001); Slavíček and Martínez (2010) We use an active space of 4 electrons in 15 orbitals. The space contains the full (four orbital) $\pi$ space, a full set of $n=3$ Rydberg orbitals (s, p, and d), a $\sigma^{*}$ orbital, and a 4s Rydberg orbital. Figure 1 depicts the active space molecular orbitals (MOs) calculated at the S0 minimum. (We define the C2 axis of rotation is the z-axis, and the molecular ring is in the yz plane.) We use the cc-pVDZ basis, extended with Rydberg basis functions positioned at the molecular center of mass (COM). These were obtained by a contraction of a set of 8s8p8d diffuse functionsKaufmann et al. (1989) with exponents optimized using the methodology of Roos et al.Roos et al. (1996) in the OpenMolcas software package.Li Manni et al. (2023) The FOMO temperature parameter is set to 0.15 a.u. This value was empirically chosen because the resulting orbitals produced excitation energies in reasonable agreement with experiment, as detailed below.
@@
To analyze the accuracy of the FOMO-CASCI excitation energies, we compare the calculated vertical excitation energies with the high-resolution absorption spectrum from Holland et al.Holland et al. (2014) in Table 1. (Transition dipole moments are given in Table S1.) The absorption spectrum contains two broad peaks centered at 5.5 and 7.0 eV. The first one is dominantly composed of valence-valence transitions (to $\pi_{1}^{*}$ and $\sigma^{*}$ states) with the ${\pi_{1}\xrightarrow{}\mathrm{3s}}$ transition on its high-energy tail. The second one then arises from a wide range of Rydberg excitations and a pair of $\pi$ ${}$ $\pi_{2}^{*}$ states. Looking at the FOMO-CASCI results, the method faithfully captures the Rydberg excitations to within a few tenths of an eV. However, FOMO-CASCI predicts the valence transitions to be  2 eV higher than observed experimentally. This is due to profound dynamic electron correlation effects which are observed mainly for the valence states, a common and well-known error in CASSCF and CASCI results, discussed above. Despite this error in excitation energy, the MO shapes, relative transition dipole moments, and shapes of the PESs are expected to be captured adequately. We will address the impact of the overestimation of the valence states on dynamical results below.

### v2

@@ I Introduction
Electronic coherences can be leveraged to control light-driven dynamics in both naturalDuan et al. (2017); Maiuri et al. (2018) and syntheticMerritt, Jacquemin, and Vacher (2021); Lépine, Ivanov, and Vrakking (2014) molecules. In an isolated atom, coherences can survive for hundreds of picoseconds or more.Tenwolde et al. (1989) Yet in molecules with more than a few atoms, the loss of electronic coherence can be ultrafast, with the vibrations of the molecule acting as a bath that kills coherence.Schwartz et al. (1996) With an eye towards control, it would be extremely desirable to understand the factors determining the rate of electronic decoherence in excited molecules.
@@
One can clearly see that a description of the vibrational wave packets on multiple electronic states is central to a correct quantitative description of electronic coherence and decoherence. Thus methods based on a fully quantum treatment of nuclear dynamics are a natural fit, providing an accurate description of coherence, albeit at a high computational cost. In fact, exact quantum simulations of model systems have shed light on numerous general features of quantum decoherence in molecules, and also provide useful benchmarks for more approximate methods.Gu and Franco (2018); Hu, Gu, and Franco (2018) When the goal is to simulate the dynamics of a particular molecule, multi-configuration time-dependent Hartree (MCTDH)Csehi et al. (2020); Dey, Kuleff, and Worth (2022); Arnold, Vendrell, and Santra (2017) can provide an extremely accurate description of the dynamics of vibrational wave packets on multiple electronic states, but requires a fitted PES, often in reduced dimensionality to keep computational costs down. Methods based on Gaussian wave packetsVacher et al. (2017); Jenkins et al. (2018); Tempelaar and Reichman (2018); Tran et al. (2020); Scheidegger, Vaníček, and Golubev (2022) provide a lower cost description in full dimensionality. But they come with their own complexities, owing to the need for multiple trajectory basis functions evolving on different PESs to interact with one another. Classical Ehrenfest simulations can provide useful insights into coherences on short timescales,Jenkins et al. (2016); Franco and Brumer (2012); Gu and Franco (2017) but Ehrenfest does not incorporate decoherence and, therefore, cannot discern between long- and short-lived coherences. This leads to well-known qualitative errors on longer timescales.Tully (1990); Cotton and Miller (2013) Incorporating decoherence effects into Ehrenfest effectively solves the problem for two-state systems,Zhu et al. (2004); Bedard-Hearn, Larsen, and Schwartz (2005) but properly treating long-lived coherences in systems with many electronic states is not straightforward.Esch and Levine (2020a) Thus, an on-the-fly classical trajectory-based approach capable of accurately describing both long-lived coherences and ultrafast decoherence has been elusive.
@@
To fill this gap, we have introduced the Ehrenfest with collapse-to-a-block (TAB) nonadiabatic molecular dynamics approach,Esch and Levine (2020a, 2021) and in this work we apply it to simulate the long-lived coherence observed in excited thiophene. The TAB method has recently been combined with an ab initio treatment of the electronic structure for the first time.Suchan et al. (2024) In the present work, we demonstrate TAB to be a computationally lightweight alternative to fully quantum methods, capable of discerning between long- and short-lived coherences. Ab initio TAB simulates the system on the fly in its full dimensionality and does not require fitted PESs. Used in conjunction with time-dependent configuration interaction (TDCI)Peng, Fales, and Levine (2018); Durden and Levine (2022) calculations of multiphoton excitation by the pump pulse, we are able to explicitly simulate the dynamics observed in the Kaufman experiment. With detailed simulations of coupled electron-nuclear dynamics in hand, we definitively assign the observed coherences to specific states and gain understanding of the physics underlying its long lifetime.
@@ II.1 Nonadiabatic dynamics: TAB method
Here, we highlight the key features of the method relevant to the present results. The TAB method is based on the Ehrenfest molecular dynamics method.Miller and McCurdy (1978); Meyera) and Miller (1979); Micha (1983); Isborn, Li, and Tully (2007) Ehrenfest is an independent trajectory method where the nuclei are treated classically, and the equation of motion is
$\frac{d}{dt}\left[\sum_{\mu}\frac{P_{\mu}^{2}}{2m_{\mu}}+\bra{\psi(t)}\hat{H}_{\mathrm{e}}(\bm{R}(t))\ket{\psi(t)}_{\bm{r}}\right]=0.$ (5)
Here $P_{\mu}$ is the $\mu$th element of the classical nuclear momentum, $m_{\mu}$ is the associated mass, and $\hat{H}_{\mathrm{e}}$ is the electronic Hamiltonian. The classical nuclear coordinates are denoted $\bm{R}(t)$, and the subscript $\bm{r}$ indicates integration over on the electronic coordinates. The electronic wave function is then evolved by solving the time-dependent electronic Schrödinger equation:
@@
In the Ehrenfest method, there is no consideration of decoherence due to the loss of nuclear overlap (mechanism c) above).Kirrander and Vacher (2020) Instead, trajectories evolve along the mean-field PES, which can result in unphysical outcomes.Tully (1990); Cotton and Miller (2013) We remedy these issues by incorporating the collapse-to-a-block (TAB) decoherence correction scheme.Esch and Levine (2020a, 2021, b)
The specific version of the TAB collapse algorithm used here is described in detail ref 36, where this correction is referred to as TAB-w1. The TAB correction assumes Gaussian decay of pairwise electronic coherences with a rate governed by the difference of the gradients of associated PESs, as originally derived by Rossky and Bittner:
@@
where $\eta$ indexes vibrational degrees of freedom, $F_{i,\eta}$ is the corresponding force matrix element on state $i$, and $\alpha_{\eta}$ is a parameter that represents the width of a Gaussian vibrational wave packet. This expression has a very intuitive interpretation; the more different the slopes of the PESs of two states, the faster their vibrational overlap, $\chi^{\prime}_{1}(\bm{R},t)\chi^{\prime*}_{2}(\bm{R},t)$, will decay. This and similar approximations have been used for years to correct the nonadiabatic transition probabilities incorporated into simulations of nonadiabatic molecular dynamics.Zhu et al. (2004); Bedard-Hearn, Larsen, and Schwartz (2005); Ouyang, Dou, and Subotnik (2015); Min, Agostini, and Gross (2015); Subotnik et al. (2016); Wang, Akimov, and Prezhdo (2016); Martens (2019); Vindel-Zandbergen et al. (2021) The novel feature of TAB is that each state pair is assigned an independent decoherence time, allowing for an accurate treatment of decoherence in systems with more than two electronic states.Esch and Levine (2020a)
@@
A complete definition of $w_{ij}(\tau_{ij},\dot{\rho}_{ii},\dot{\rho}_{jj})$ may be found in ref 36. The diagonal elements of the target density matrix are not scaled,
@@
Finally, we stochastically decide whether the wavefunction will collapse into one of pure states represented by these blocks. A random number between 0 and 1 is selected and compared to the probabilities, $\{P_{a}^{\mathrm{block}}\}$, to determine which state the electronic subsystem collapses into.
After collapse, the entire momentum vector is scaled to ensure conservation of energy with a factor of $\sqrt{1-\frac{E_{\mathrm{pot}}^{\mathrm{new}}-E_{\mathrm{pot}}^{\mathrm{old}}}{E_{\mathrm{kin}}}}$, adjusting the kinetic energy $E_{\mathrm{kin}}$ by the difference of TDCI electronic energies $E_{\mathrm{pot}}$ after and before the collapse. This block-wise algorithm allows TAB to correctly describe the decay of coherences between all pairs of states—even when many states are present, pairs of states with parallel PESs maintain coherence indefinitely, while very non-parallel pairs lose coherence rapidly.Esch and Levine (2020a) Like the parent Ehrenfest method, a single TAB trajectory is not meaningful. However, averaging over an ensemble yields an accurate treatment of decoherence, with subset of trajectories ultimately collapsing to individual states or sets of parallel states after a physically meaningful decoherence time.
In this work, we use a fully ab initio method to propagate the electronic wave function: time-dependent configuration interaction.Peng, Fales, and Levine (2018) In TDCI, the wave function is expressed as a linear combination of Slater determinants, $\Phi_{I}$, with time-dependent expansion coefficients $C_{I}(t)$,
@@
$i\dot{\bm{C}}(t)={\bm{H}_{\mathrm{e}}}(t){\bf{C}}(t)$ (12)
and solve the equation numerically using the symplectic split operator integrator.Blanes, Casas, and Murua (2006) In this work, we select a complete active space (TD-CASCI) ansatz. Our GPU-accelerate implementation is in the TeraChem electronic structure software package.Seritan et al. (2021, 2020); Durden and Levine (2022); Fales and Levine (2015); Peng, Fales, and Levine (2018); Hohenstein et al. (2015)
To derive an expression for the forces experienced by the classical nuclei, we must note that the time-dependence of the electronic Hamiltonian matrix arises not only from an explicit dependence of the Hamiltonian operator on $\bm{R}(t)$, but also a dependence of the many-electron basis on the position-dependent orbitals, $\bm{X}(\bm{R}(t))$. Then, starting from Eq. 5 we can derive the expression for the $\mu$th element of the force,
$F_{\mu}=-\bm{C}^{\dagger}\frac{\partial\bm{H}_{\mathrm{e}}}{\partial R_{\mu}}\bm{C}-\left(\bm{C}^{\dagger}\frac{\partial\bm{H}_{\mathrm{e}}}{\partial\bm{X}}\bm{C}\right)\frac{\partial\bm{X}}{\partial R_{\mu}}.$ (13)
All terms are computed using existing time-independent gradient code, with only minor modifications.Hohenstein et al. (2015)
We have previously applied ab initio TAB to predict the ultrafast electronic diffraction spectrum of the photodissociation of cyclobutanoneSuchan et al. (2024) in the context of a community prediction challenge.Miao, Diemer, and Mitrić (2024); Martín Santa Daría et al. (2024) Our prediction, though slightly too fast, is in good agreement with experiment. [^1] Preliminary experimental data communicated by Thomas Wolf. However, that problem involved a small number of electronic states and only short-lived coherences. The present work represents the first application of TAB to describe long-lived electronic coherences in a system with many electronic states, the class of problems for which it is almost uniquely well suited.
@@ II.2 Thiophene’s electronic structure
Despite its apparent simplicity, previous theoretical studies reveal that accurately describing the electronic excited states of thiophene (C4H4S) can be quite challenging. The HOMO ($\pi_{1}$) and HOMO-1 ($\pi_{2}$) orbitals, from which the low-lying excitations originate, have $\pi$ character. The excitations lead not only to antibonding valence orbitals ($\pi^{*}$ and $\sigma^{*}$), but also to a set of low-lying Rydberg orbitals. This puts additional demands on the atomic basis set and wave function expansion. The difference in electronic correlation energy between the valence and Rydberg states is quite large, as demonstrated in many past studies based on configuration interactionPalmer, Walker, and Guest (1999), complete active space self-consistent field (CASSCF)Serrano-Andrés et al. (1993); Heller, Joswig, and Seifert (2021); Pastore, Angeli, and Cimiraglia (2007), multireference perturbation theory Kaufman et al. (2022); Parkes and Worth (2024), coupled cluster theory Pastore, Angeli, and Cimiraglia (2007); Holland et al. (2014), and other methods.Huang et al. (2008); Salzmann et al. (2008); Kleinschmidt, Tatchen, and Marian (2002); Jones et al. (2019)
For computational feasibility, we must choose a relatively inexpensive electronic structure method for use in conjunction with TAB. To this end, we choose the time-dependent (TD-) floating occupation molecular orbitals (FOMO-) complete active space configuration interaction (CASCI) approach.Granucci, Persico, and Toniolo (2001); Slavíček and Martínez (2010) We use an active space of 4 electrons in 15 orbitals. The space contains the full (four orbital) $\pi$ space, a full set of $n=3$ Rydberg orbitals (s, p, and d), a $\sigma^{*}$ orbital, and a 4s Rydberg orbital. Figure 1 depicts the active space molecular orbitals (MOs) calculated at the S0 minimum. (We define the C2 axis of rotation is the z-axis, and the molecular ring is in the yz plane.) We use the cc-pVDZ basis, extended with Rydberg basis functions positioned at the molecular center of mass (COM). These were obtained by a contraction of a set of 8s8p8d diffuse functionsKaufmann, Baumeister, and Jungen (1989) with exponents optimized using the methodology of Roos et al.Roos et al. (1996) in the OpenMolcas software package.Li Manni et al. (2023) The FOMO temperature parameter is set to 0.15 a.u. This value was empirically chosen because the resulting orbitals produced excitation energies in reasonable agreement with experiment, as detailed below.
@@
To analyze the accuracy of the FOMO-CASCI excitation energies, we compare the calculated vertical excitation energies with the high-resolution absorption spectrum from Holland et al.Holland et al. (2014) in Table 1. (Transition dipole moments are given in Table S1.) The absorption spectrum contains two broad peaks centered at 5.5 and 7.0 eV. The first one is dominantly composed of valence-valence transitions (to $\pi_{1}^{*}$ and $\sigma^{*}$ states – lower part of Table 1), with the ${\pi_{1}\xrightarrow{}\mathrm{3s}}$ transition on its high-energy tail. The second one then arises from a wide range of Rydberg excitations and a pair of $\pi$ ${}$ $\pi_{2}^{*}$ states. Looking at the FOMO-CASCI results, the method faithfully captures the Rydberg excitations to within a few tenths of an eV. However, FOMO-CASCI predicts the valence transitions to be  2 eV higher than observed experimentally. This is due to profound dynamic electron correlation effects which are observed mainly for the valence states, a common and well-known error in CASSCF and CASCI results, discussed above. Despite this error in excitation energy, the MO shapes, relative transition dipole moments, and shapes of the PESs are expected to be captured adequately. We will address the impact of the overestimation of the valence states on dynamical results below.